​Suppose you bought a bond with an annual coupon rate of 5.5 percent one year ago for $1,017

Suppose you bought a bond with an annual coupon rate of 5.5 percent one year ago for $1,017. The bond sells for $1,041 today. 


a. Assuming a $1,000 face value, what was your total dollar return on this investment over the past year? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) 

b. What was your total nominal rate of return on this investment over the past year? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

c. If the inflation rate last year was 3 percent, what was your total real rate of return on this investment? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

a. Total dollar return = _______ 

b. Total nominal rate of return = _______ %

c. Total real rate of return = _______ %
Answer #1

Bond purchase price=   $1,017  
Current or sale price=   $1,041  
one year coupon received = face value*coupon rate      
1000*5.5%=   55  


(A) Dollar return formula = Closing price - purchase price+coupon received)l      
1041-1017+55      
$79      
so total dollar return is    $79  
      
(B)       
Nominal Rate of return formula = (Sale price - purchase price+coupon received)/Purchase price      
(1041-1017+55)/1017      
      
0.07767945   or 7.77%  
      
So, Total nominal rate of return on bond is 7.77%       
      
(C) Real rate if return = ((1+Nominal Rate of return)/(1+inflation rate))-1      
((1+0.07767945)/(1+3%))-1      
0.0463   or 4.63%  
So real rate of return earned is   4.63%  
      

Please thumbs up
Answer #2

a.

Dollar Return = (1,041 - 1,017 + 55) = $79

b.

Nominal Rate = 79/1,017 = 7.77%

c.

Real Rate = (1.077/1.03) - 1 = 4.56%
Answer #3

1.

=(Face value*Coupon rate+Price today-Price one year ago)

=(1000*5.5%+1041-1017)=79

2.

=Return/Purchase Price

=79/1017=7.77%

3.

=(1+7.77%)/(1+3%)-1

=4.63%